OpenStudy (anonymous):

How do you determine a maximum and minimum value without graphing. I was given an equation..could that help?

OpenStudy (anonymous):

Calculus is the easiest way, though you said in an earlier question that you haven't taken it yet. Perhaps give your equation, since there are non-calculus techniques that can be used for certain types of equations, like sinusoidal curves, that I could show you.

OpenStudy (anonymous):

alright my equation follows: \[P(t) = 5000\sin(30(t-7)) + 8000 where t is the number of months after New Years

OpenStudy (anonymous):

ignore that slash after 5000

OpenStudy (anonymous):

Ah a quick note, if you want openstudy to register what you write as an equation, you also have to end it with \]. And you can leave it as \sin, rather than sin, it makes it look a little nicer. To answer your question, do you know the range of values that sine can take?

OpenStudy (anonymous):

Observe using \sin \[\sin x\] versus just sin \[sin x\]

OpenStudy (anonymous):

\[P(t) = 5000\sin (30(t-7)) + 8000\]

OpenStudy (anonymous):

what do you mean by the range of values sin can take?

OpenStudy (anonymous):

Okay, if you consider \[y = \sin x\] where x can be any real number, what are all possible values that y can be?

OpenStudy (anonymous):

y = -1, 0, 1

OpenStudy (anonymous):

Even more than that, sine can take on ALL values between -1 and 1 (inclusive). If you consider what a general sine curve looks like, you can see this for yourself. Now, since sin x can be anything from -1 to 1 and will ALWAYS take on those values, if you want to find the maxima and minima of a sinusoidal curve, say \[f(x)=A\sin(Bx-C)+D\] Then the smallest value sin can be is -1 and the largest it can be is 1. Thus, if you plug in 1 for sin, you will get the maxima: \[f(x)=A(1)+D=A+D\] and plugging in -1 for sin gives the minima: \[f(x)=A(-1)+D=-A+D\] This works for all sine and cosine curves.

OpenStudy (anonymous):

thankkksss :) but for the value (t) what do i plug in 12 months?

OpenStudy (anonymous):

Well, if you want the maximum of this curve, you would just need to find points where \[\sin(30(t-7))=1\] This will equal 1 an infinite number of times and a little trig will show that it equals 1 if \[30(t-7)=\frac{\pi}{2}+2\pi k\] for ANY integer value of k (..., -3, -2, -1, 0, 1, 2, 3, ...) If you don't care which maximum point you have, just set k=0. Then you have \[30(t-7)=\frac{\pi}{2}\] which you can easily solve for t. Unfortunately, as you will often see with trigonometry, the answer isn't a simple number like 12.

OpenStudy (anonymous):

you seriously explain things far better than my math teacher thanks dude (: well i wont be focusing on pi till next year although i figured the answers to this question using the steps you showed. Thanks again

OpenStudy (anonymous):

No problem :) I love teaching and I'm always glad to help when people need it.

OpenStudy (zarkon):

I'd be surprised if he was working in radians...I'm betting degrees.

OpenStudy (anonymous):

Ah yes, that is true, if you replace the 30 with pi/6, then you get 10 months. I forgot that you are using degrees in your answers and combined your degrees with my radians.

OpenStudy (anonymous):

Aha yeah i got 10 months for my maxima and for my minima i got 4 months. So far so good. I got 2 more questions to go until starting on my chemistry.